# Morphology of 32 Repeating Fast Radio Burst Sources at Microsecond Time Scales with CHIME/FRB

### v1

@@ 4.1.5 Extragalactic DM & Scattering Correlation
Only 9 of the repeating FRB sources in our sample have at least one burst with a measured scattering timescale. Thus, the sample for this correlation is extremely limited. Nonetheless, the Spearman p-value using only the sources with measured scattering timescales is 0.63222Given the low number of sources, and the low significance, we do not perform the same MC employed for previous correlations.. We also run a censored Kendall test using the cenken function in R’s NADA package(Lopaka, 2020; R Core Team, 2020). The censored Kendall test allows us to account for upper limits on our scattering timescales (also referred to as left censored data). The Kendall value is 0.1, suggesting little to no correlation. The lack of a correlation supports the notion that a significant portion of the extragalactic DM does not arise from the same location as that of the scattering. Our results are similar to those of Sand et al. (2024).
@@
We compare various morphological features between our sample of repeating FRBs and the thus-far non-repeating FRB sample presented by Sand et al. (2024). A significant advantage to comparing these two populations rather than including FRBs from instruments other than CHIME/FRB is that both suffer from the same instrumental biases, aside from their detection thresholds. The difference in detection thresholds is the most relevant for comparisons of the spectral fluences (Fν) of the bursts. Thus, for all fluence comparisons, we only include bursts for which the real-time detection S/N333While the S/N threshold for non-repeaters is 12, we choose 20 to be well above the completeness level. . For other tests, we place cuts on the time resolution at which the burst is studied. Thus, this is approximately equivalent to a S/N cut, as only bursts with sufficient S/N can be studied at high time resolutions. We compare the bandwidth, total burst duration, number of burst components, sub-burst widths, sub-burst waiting times, and fluences between the two samples in Figures 2 - 6.
To assess the difference between the two populations, we use both the Anderson-Darling (AD; Anderson & Darling, 1954) and Kolmogorov-Smirnov (KS; KST, 2008) test statistics. Similar to in Section 4.1, we consider two samples to be distinct if the AD and KS test p-values are both . Results from all of our comparisons are listed in Table 1. We include the total number of repeating and non-repeating FRBs (or the number of bursts) used per comparison after applying various cuts in Table 1. We perform the comparisons using each individual burst from a given source, rather than the average of that quantity for the source. Thus, repeaters with a significant number of bursts will more heavily contribute to the given distributions.
@@
| Bandwidth | N/A | 118 | 125 | ||
| Duration | N/A | 118 | 125 | ||
| Num. Components | t ms | 20 | 79 | 0.25 | 0.99 |
@@
| Sub-burst Width | ts | 18eeNumber of individual sub-bursts rather than number of detected bursts. | 164eeNumber of individual sub-bursts rather than number of detected bursts. | 0.003 | 0.003 |
| Duration-Normalized Sub-burst Width | t ms | 89eeNumber of individual sub-bursts rather than number of detected bursts. | 210eeNumber of individual sub-bursts rather than number of detected bursts. | 0.24 | 0.15 |
| Duration-Normalized Sub-burst Width | ts | 36eeNumber of individual sub-bursts rather than number of detected bursts. | 190eeNumber of individual sub-bursts rather than number of detected bursts. | 0.015 | 0.0031 |
| Duration-Normalized Sub-burst Width | ts | 18eeNumber of individual sub-bursts rather than number of detected bursts. | 164eeNumber of individual sub-bursts rather than number of detected bursts. | 0.0032 | 0.0013 |
| Wait time | t ms | 48ffNumber of arrival time differences. Only applicable for multi-component bursts. | 113ffNumber of arrival time differences. Only applicable for multi-component bursts. | ||
| Wait time | ts | 15ffNumber of arrival time differences. Only applicable for multi-component bursts. | 108ffNumber of arrival time differences. Only applicable for multi-component bursts. | ||
| Wait time | ts | 6ffNumber of arrival time differences. Only applicable for multi-component bursts. | 98ffNumber of arrival time differences. Only applicable for multi-component bursts. | 0.01 | 0.01 |
| Duration-Normalized Wait time | t ms | 48ffNumber of arrival time differences. Only applicable for multi-component bursts. | 113ffNumber of arrival time differences. Only applicable for multi-component bursts. | 0.25 | 0.96 |
| Duration-Normalized Wait time | ts | 15ffNumber of arrival time differences. Only applicable for multi-component bursts. | 108ffNumber of arrival time differences. Only applicable for multi-component bursts. | 0.013 | 0.14 |
| Duration-Normalized Wait time | ts | 6 | 98 | 0.006 | 0.031 |
@@
| Band-limited FluenceddSpectral fluence averaged over the spectral extent of the burst. We limit the sample to bursts whose detection S/N is greater than 20. | S/N | 35 | 48 | 0.25 | 0.48 |
@@
CHIME/FRB Collaboration et al. (2023a) found a significant difference between the DM distributions of repeating and non-repeating FRBs. This DM distribution difference persisted as the S/N threshold of their sample was lowered, as well as when extragalactic DM instead of total DM was considered. They found an extragalactic DM difference of 161 55 pc cm-3 between repeaters and non-repeaters detected above a S/N threshold of 12, with repeaters having on average lower extragalactic DMs. While our analysis has been able to refine the DM of repeating FRBs, the improvement has been on the order of 1 pc cm-3 or less. Hence, we do not expect a significant change from the results of CHIME/FRB Collaboration et al. (2023a) regarding the DM distribution.
@@
To compare the two populations, we use the Kaplan-Meier method, a part of the python package lifelines, and the Peto-Peto test statistic (Peto & Peto, 1972). These methods give significant advantages over the AD and KS tests as data can be labeled as either censored or uncensored. In the case of scattering timescales, the censored data are those for which we have only upper limits, while the uncensored data are the measured scattering timescales. In the right panel of Figure 2, we show the cumulative distribution function (CDF) of the repeating and non-repeating scattering distributions using the Kaplan-Meier method. The Peto-Peto test statistic for the two distributions is . Thus, given the current sample sizes, there is no evidence for a difference in the scattering timescales of repeating and thus-far non-repeating FRBs. However, we caution that our repeating FRB sample is small, and those with measured scattering timescales even smaller. Additional repeating FRBs with measured scattering timescales could still reveal a difference between the two populations.
@@
In an effort to systematically search for other apparent differences in emission phenomenology between repeaters and non-repeaters, we compare the number of sub-bursts per event444Note this is not per source, but rather per individual burst from a given source., the width of individual sub-bursts, and the waiting time (defined as the difference in the time of arrival of individual sub-bursts) between the two populations. However, all of these comparisons can be biased by the time resolution at which we study the populations as additional components may be unresolved at lower resolutions. To try to address this bias, we limit the comparison to bursts (both repeating and non-repeating) studied at a time resolution 0.16 ms. However, a bias may still remain as many of the non-repeating FRBs are bright enough to be analyzed at time resolutions 20.48 s (e.g., 51 out of 125 bursts) while only 7 bursts from our repeating sample are studied at time resolutions 20.48 s. Additionally, we note that more active repeaters will influence the distributions more than less active sources, as we use every burst from each source with the required data for our comparisons.
@@
There is an inherent bias in a comparison of the spectral fluences555Unless specified otherwise, the spectral fluences referred to throughout this work will be those averaged over the full 400 to 800 MHz CHIME/FRB band. of the two populations since the triggering threshold at CHIME/FRB for the complex raw voltages is lower for repeaters than for non-repeaters. Thus, we only include bursts in our sample for which the initial detection S/N was , as this is sufficiently above the triggering threshold for both repeating and non-repeating sources. Additionally, missing channels due to RFI masking could also affect this comparison. However, given we are only comparing bursts detected by CHIME, we do not expect drastic changes in the total fraction of the band that is masked due to RFI. We remove bursts for which the top available channel was less than 800 MHz, though (e.g., those for which DM pc cm-3).
@@ 4.3 Spectral Fluence, Bandwidth, DM, and Scattering Time Variations
If FRBs are young NS embedded in turbulent environments, we might expect the local environment (probed through the DM, rotation measure, scattering timescale, and scintillation features) of these sources to change with time (Piro & Gaensler, 2018).666We note, however, that changes in the DM, rotation measure, etc. could also be caused by the relative motion of the source with respect to us and thus a changing DM does not directly imply a changing environment. Michilli et al. (2018) found extreme changes in the rotation measure (RM) of FRB 20121102A (which may be embedded in a young nebula e.g., see Marcote et al., 2017) while Hessels et al. (2019) found evidence for DM changes of 1-4 pc cm-3 over a four year period for this source. Mckinven et al. (2023b) and Ng et al. (in prep.) also both find that most repeating FRBs show subtle RM changes of order a few rad m-2, while some show significant RM variations. However, Mckinven et al. (2023b) did not perform detailed studies of the DMs of the sources in their analysis. Bethapudi et al. (2022) also found changes in the scintillation bandwidth for FRB 20180916B with time.
@@
Here, we explore whether any sources in our sample show changes in their DM and scattering timescale as a function of time. In particular, we focus on the repeating sources FRB 20191106C, FRB 20200929C, FRB 20181119A, FRB 20190208A, FRB 20190303A, and FRB 20190417A. These sources are chosen as they each have six or more bursts with the complex raw voltages recorded777A minimum of six bursts was chosen in order to have a sufficiently large sample of burst property measurements while also enabling examination of multiple sources from our sample.. In addition to DM and scattering timescales, we also explore whether the spectral fluence of a burst impacts (or is impacted by) any other burst property and whether there is a systematic change in the emitting bandwidth of these sources (largely motivated by FRB 20180916B, whose activity window is frequency dependent e.g., see Pastor-Marazuela et al., 2021; Pleunis et al., 2021b; Bethapudi et al., 2022). To quantify changes in a given parameter, we perform a weighted least-squares fit assuming a linear function. We do not include bursts for which we either calculate an upper limit on the scattering timescale, or for which we do not report a DM uncertainty.
@@
In Figures 7 - 12, we show the fluence, DM, scattering timescale (if applicable888Only shown if we can measure a scattering timescale for at least one of the bursts from the source.), and frequency range for FRBs 20181119A, 20190208A, 20190303A, 20190417A, 20191106C, and 20200929C, respectively. The best-fit slopes for spectral fluence, DM, and scattering timescale (along with their 1 uncertainties) are listed in Table 2. Except for the temporal variation of the spectral fluence of FRB 20181119A, all other derived slopes are either not statistically significant or are due to underestimated measurement uncertainties. We do not perform the same comparison for bandwidth, as the central frequency is highly dependent on the CHIME/FRB beam response which we have not corrected for in this work. Thus, for bandwidth, only by-eye trends are reported999The actual bandwidths are still calculated using the fitburst model assuming a running power law..
@@ 4.3.1 Spectral Fluence Variations over Time
FRB 20181119A is the only source for which there is a significant (slope greater than the 3 uncertainties) change in the spectral fluence (over the entire 400-MHz CHIME/FRB band) with time (see Figure 7). The best fit slope is Jy ms yr-1.
@@ 4.3.3 DM Variations over Time
Other than FRB 20181119A (Figure 7), none of the sources in our sample have a significant secular increase or decrease in DM as a function of time. While the trend for FRB 20181119A is significant, it is based on only three data points, and does not include the final two measurements that were estimated by-eye. Additionally, we manually de-disperse all bursts from this source to the weighted-average DM and all bursts still appear fully de-dispersed. Thus, the variation for this source is likely not astrophysical and is instead due to difficulty in determining the DMs of the bursts as many have low S/Ns without obvious sub-structure.
@@ 5.1 Local Environments
As discussed in Section 4.2, we do not find a statistical difference in the scattering timescales between repeating and non-repeating FRBs. Pandhi et al. (2024) found no significant difference in the linear polarization fraction or the extragalactic RM when comparing these two populations. They did not, however, find any thus-far non-repeaters with extreme RMs, in contrast with the extreme RMs seen in the repeating FRBs 20121102 and 20190520B (Michilli et al., 2018; Anna-Thomas et al., 2023). This could be explained by intra-channel depolarization in the CHIME/FRB band, though. Nonetheless, taken together, these findings suggest that repeating and non-repeating FRBs share similar local environments. This supports the notion that the differences observed in the DM distributions by CHIME/FRB Collaboration et al. (2023a) are unlikely to be driven by local environmental factors. Nonetheless, the small number of known repeating FRBs compared to non-repeaters suggests that these conclusions should be viewed cautiously, as additional data may reveal more subtle differences.

### v2

@@ 4.1.5 Extragalactic DM & Scattering Correlation
Only 9 of the repeating FRB sources in our sample have at least one burst with a measured scattering timescale. Thus, the sample for this correlation is extremely limited. Nonetheless, the Spearman p-value using only the sources with measured scattering timescales is 0.63333Given the low number of sources, and the low significance, we do not perform the same MC employed for previous correlations.. We also run a censored Kendall test using the cenken function in R’s NADA package(Lopaka, 2020; R Core Team, 2020). The censored Kendall test allows us to account for upper limits on our scattering timescales (also referred to as left censored data). The Kendall value is 0.1, suggesting little to no correlation. The lack of a correlation supports the notion that a significant portion of the extragalactic DM does not arise from the same location as that of the scattering. Our results are similar to those of Sand et al. (2024).
@@
We compare various morphological features between our sample of repeating FRBs and the thus-far non-repeating FRB sample presented by Sand et al. (2024). A significant advantage to comparing these two populations rather than including FRBs from instruments other than CHIME/FRB is that both suffer from the same instrumental biases, aside from their detection thresholds. The difference in detection thresholds is the most relevant for comparisons of the spectral fluences (Fν) of the bursts. Thus, for all fluence comparisons, we only include bursts for which the real-time detection S/N444While the S/N threshold for non-repeaters is 12, we choose 20 to be well above the completeness level. . For other tests, we place cuts on the time resolution at which the burst is studied. Thus, this is approximately equivalent to a S/N cut, as only bursts with sufficient S/N can be studied at high time resolutions. We compare the bandwidth, total burst duration, number of burst components, sub-burst widths, sub-burst waiting times, and fluences between the two samples in Figures 2 - 6.
To assess the difference between the two populations, we use both the Anderson-Darling (AD; Anderson & Darling, 1954) and Kolmogorov-Smirnov (KS; KST, 2008) test statistics. Similar to in Section 4.1, we consider two samples to be distinct if the AD and KS test p-values are both . Results from all of our comparisons are listed in Table 1. We include the total number of bursts from repeating and non-repeating FRBs used per comparison after applying various cuts in Table 1. We perform the comparisons using each individual burst from a given source, rather than the average of that quantity for the source. Thus, repeaters with a significant number of bursts will more heavily contribute to the given distributions.
@@
Bandwidth | N/A | 124 | 125 | ||
Duration | N/A | 124 | 125 | ||
Num. Components | t ms | 20 | 79 | 0.3 | 0.7 |
@@
Sub-burst Width | ts | 18eeNumber of individual sub-bursts rather than number of detected bursts. | 164eeNumber of individual sub-bursts rather than number of detected bursts. | 0.008 | 0.007 |
Duration-Normalized Sub-burst Width | t ms | 89eeNumber of individual sub-bursts rather than number of detected bursts. | 210eeNumber of individual sub-bursts rather than number of detected bursts. | 0.3 | 0.4 |
Duration-Normalized Sub-burst Width | ts | 36eeNumber of individual sub-bursts rather than number of detected bursts. | 190eeNumber of individual sub-bursts rather than number of detected bursts. | 0.02 | 0.009 |
Duration-Normalized Sub-burst Width | ts | 18eeNumber of individual sub-bursts rather than number of detected bursts. | 164eeNumber of individual sub-bursts rather than number of detected bursts. | 0.008 | 0.007 |
Wait time | t ms | 48ffThe number of unique waiting times calculated as the differences between arrival times. Only applicable for multi-component bursts. | 113ffThe number of unique waiting times calculated as the differences between arrival times. Only applicable for multi-component bursts. | ||
Wait time | ts | 15ffThe number of unique waiting times calculated as the differences between arrival times. Only applicable for multi-component bursts. | 108ffThe number of unique waiting times calculated as the differences between arrival times. Only applicable for multi-component bursts. | ||
Wait time | ts | 6ffThe number of unique waiting times calculated as the differences between arrival times. Only applicable for multi-component bursts. | 98ffThe number of unique waiting times calculated as the differences between arrival times. Only applicable for multi-component bursts. | 0.02 | 0.02 |
Duration-Normalized Wait time | t ms | 48ffThe number of unique waiting times calculated as the differences between arrival times. Only applicable for multi-component bursts. | 113ffThe number of unique waiting times calculated as the differences between arrival times. Only applicable for multi-component bursts. | 0.25 | 0.98 |
Duration-Normalized Wait time | ts | 15ffThe number of unique waiting times calculated as the differences between arrival times. Only applicable for multi-component bursts. | 108ffThe number of unique waiting times calculated as the differences between arrival times. Only applicable for multi-component bursts. | 0.01 | 0.06 |
Duration-Normalized Wait time | ts | 6 | 98 | 0.005 | 0.02 |
@@
Band-limited FluenceddSpectral fluence averaged over the spectral extent of the burst. We limit the sample to bursts whose detection S/N is greater than 20. | S/N | 35 | 48 | 0.3 | 0.5 |
@@
CHIME/FRB Collaboration et al. (2023a) found a significant difference between the DM distributions of repeating and non-repeating FRBs. This DM distribution difference persisted as the S/N threshold of their sample was lowered, as well as when extragalactic DM instead of total DM was considered. They found an extragalactic DM difference of 161 55 pc cm-3 between the mean DM of repeaters and non-repeaters detected above a S/N threshold of 12, with repeaters having on average lower extragalactic DMs. While our analysis has been able to refine the DM of repeating FRBs, the improvement has been on the order of 1 pc cm-3 or less. Hence, we do not expect a significant change from the results of CHIME/FRB Collaboration et al. (2023a) regarding the DM distribution.
@@
To compare the two populations, we use the Kaplan-Meier method, a part of the python package lifelines, and the Peto-Peto test statistic (Peto & Peto, 1972). These methods give significant advantages over the AD and KS tests as data can be labeled as either censored or uncensored. In the case of scattering timescales, the censored data are those for which we have only upper limits, while the uncensored data are the measured scattering timescales. In the right panel of Figure 2, we show the cumulative distribution function (CDF) of the repeating and non-repeating scattering distributions using the Kaplan-Meier method. The Peto-Peto test statistic for the two distributions is (significantly larger than our significance criteria of ). Thus, given the current sample sizes, there is no evidence for a difference in the scattering timescales of repeating and thus-far non-repeating FRBs. However, we caution that our repeating FRB sample is small, and those with measured scattering timescales even smaller. Additional repeating FRBs with measured scattering timescales could still reveal a difference between the two populations.
@@
In an effort to systematically search for other apparent differences in emission phenomenology between repeaters and non-repeaters, we compare the number of sub-bursts per event555Note this is not per source, but rather per individual burst from a given source., the width of individual sub-bursts, and the waiting time (defined as the difference in the time of arrival of individual sub-bursts) between the two populations. However, all of these comparisons can be biased by the time resolution at which we study the populations as additional components may be unresolved at lower resolutions. To try to address this bias, we limit the comparison to bursts (both repeating and non-repeating) studied at a time resolution 0.16 ms. However, a bias may still remain as many of the non-repeating FRBs are bright enough to be analyzed at time resolutions 20.48 s (e.g., 51 out of 125 bursts) while only 7 bursts from our repeating sample are studied at time resolutions 20.48 s. Additionally, we note that more active repeaters will influence the distributions more than less active sources, as we use every burst from each source with the required data for our comparisons.
@@
There is an inherent bias in a comparison of the spectral fluences666Unless specified otherwise, the spectral fluences referred to throughout this work will be those averaged over the full 400 to 800 MHz CHIME/FRB band. of the two populations since the triggering threshold at CHIME/FRB for the complex raw voltages is lower for repeaters than for non-repeaters. Thus, we only include bursts in our sample for which the initial detection S/N was , as this is sufficiently above the triggering threshold for both repeating and non-repeating sources. Additionally, missing channels due to RFI masking could also affect this comparison. However, given we are only comparing bursts detected by CHIME, we do not expect drastic changes in the total fraction of the band that is masked due to RFI. We remove bursts for which the top available channel was less than 800 MHz, though (e.g., those for which DM pc cm-3) as the fluence may biased high if we assume a negative spectral index.
@@ 4.3 Spectral Fluence, Bandwidth, DM, and Scattering Time Variations
If FRBs are young NS embedded in turbulent environments, we might expect the local environment (probed through the DM, rotation measure, scattering timescale, and scintillation features) of these sources to change with time (Piro & Gaensler, 2018).777We note, however, that changes in the DM, rotation measure, etc. could also be caused by the relative motion of the source with respect to us and thus a changing DM does not directly imply a changing environment. Michilli et al. (2018) found extreme changes in the rotation measure (RM) of FRB 20121102A (which may be embedded in a young nebula e.g., see Marcote et al., 2017) while Hessels et al. (2019) found evidence for DM changes of 1-4 pc cm-3 over a four year period for this source. Mckinven et al. (2023b) and Ng et al. (in prep.) also both find that most repeating FRBs show subtle RM changes of order a few rad m-2, while some show significant RM variations. However, Mckinven et al. (2023b) did not perform detailed studies of the DMs of the sources in their analysis. Bethapudi et al. (2023) also found changes in the scintillation bandwidth for FRB 20180916B with time.
@@
Here, we explore whether any sources in our sample show changes in their DM and scattering timescale as a function of time. In particular, we focus on the repeating sources FRB 20191106C, FRB 20200929C, FRB 20181119A, FRB 20190208A, FRB 20190303A, and FRB 20190417A. These sources are chosen as they each have six or more bursts with the complex raw voltages recorded888A minimum of six bursts was chosen in order to have a sufficiently large sample of burst property measurements while also enabling examination of multiple sources from our sample.. In addition to DM and scattering timescales, we also explore whether the spectral fluence of a burst impacts (or is impacted by) any other burst property and whether there is a systematic change in the emitting bandwidth of these sources (largely motivated by FRB 20180916B, whose activity window is frequency dependent e.g., see Pastor-Marazuela et al., 2021; Pleunis et al., 2021b; Bethapudi et al., 2023). To quantify changes in a given parameter, we perform a weighted least-squares fit assuming a linear function. We do not include bursts for which we either calculate an upper limit on the scattering timescale, or for which we do not report a DM uncertainty.
@@
In Figures 7 - 12, we show the fluence, DM, scattering timescale (if applicable999Only shown if we can measure a scattering timescale for at least one of the bursts from the source.), and frequency range for FRBs 20181119A, 20190208A, 20190303A, 20190417A, 20191106C, and 20200929C, respectively. The best-fit slopes for spectral fluence, DM, and scattering timescale (along with their 1 uncertainties) are listed in Table 2. Except for the temporal variation of the spectral fluence of FRB 20181119A, all other derived slopes are either not statistically significant or are due to underestimated measurement uncertainties. We do not perform the same comparison for bandwidth, as the central frequency is highly dependent on the CHIME/FRB beam response which we have not corrected for in this work. Thus, for bandwidth, only by-eye trends are reported101010The actual bandwidths are still calculated using the fitburst model assuming a running power law..
@@ 4.3.1 Spectral Fluence Variations over Time
FRB 20181119A is the only source for which there is a significant (slope greater than the 3 uncertainties) change in the spectral fluence (over the entire 400-MHz CHIME/FRB band) with time (see Figure 7). The best fit slope is Jy ms yr-1. However, the reduced -squared for this fit is 25, implying a poor fit to the data.
@@ 4.3.3 DM Variations over Time
Other than FRB 20181119A (Figure 7), none of the sources in our sample have a significant secular increase or decrease in DM as a function of time. While the trend for FRB 20181119A is significant, it is based on only four data points, and does not include the final two measurements that were estimated by-eye. Additionally, the reduced chi-squared for this fit is 60, suggesting it is a very poor fit to the data. We also manually de-disperse all bursts from this source to the weighted-average DM and all bursts still appear fully de-dispersed. Thus, the variation for this source is likely not astrophysical and is instead due to difficulty in determining the DMs of the bursts as many have low S/Ns without obvious sub-structure.
@@ 5.1 Local Environments
As discussed in Section 4.2, we do not find a statistical difference in the scattering timescales between repeating and non-repeating FRBs. Pandhi et al. (2024) found no significant difference in the linear polarization fraction or the extragalactic RM when comparing these two populations. They did not, however, find any thus-far non-repeaters with extreme RMs, in contrast with the extreme RMs seen in the repeating FRBs 20121102 and 20190520B (Michilli et al., 2018; Anna-Thomas et al., 2023). This could be explained by intra-channel depolarization in the CHIME/FRB band as CHIME/FRB would have an approximately 50% decrease in its sensitivity to bursts with RMs 5000 rad m-2 at 600 MHz due to depolarization (Mckinven et al., 2021).
Nonetheless, taken together, these findings suggest that repeating and non-repeating FRBs share similar local environments. This supports the notion that the differences observed in the DM distributions by CHIME/FRB Collaboration et al. (2023a) are unlikely to be driven by local environmental factors. Nonetheless, the small number of known repeating FRBs compared to non-repeaters suggests that these conclusions should be viewed cautiously, as additional data may reveal more subtle differences.